If your concave plate with a mass of 2 kg and spinning at a velocity of 4 rad/s, what is its angular momentum?
1
Expert's answer
2020-04-17T10:19:54-0400

r is the distance from the pivot.

Angular momentum:


"M=M\omega r^2=(2)(4) r^2=(8r^2)\frac{kg\cdot m^2}{s}"